### v3

# Craw4LLM: Efficient Web Crawling for LLM Pretraining

Abstract
Web crawl is a main source of large language models’ (LLMs) pretraining data, but the majority of crawled web pages are discarded in pretraining due to low data quality. This paper presents Craw4LLM, an efficient web crawling method that explores the web graph based on the preference of LLM pretraining. Specifically, it leverages the influence of a webpage in LLM pretraining as the priority score of the web crawler’s scheduler, replacing the standard graph-connectivity-based priority. Our experiments on a web graph containing 900 million webpages from a commercial search engine’s index demonstrate the efficiency of Craw4LLM in obtaining high-quality pretraining data. With just 21% URLs crawled, LLMs pretrained on Craw4LLM data reach the same downstream performances of previous crawls, significantly reducing the crawling waste and alleviating the burdens on websites. Our code is publicly available at https://github.com/cxcscmu/Craw4LLM.
Craw4LLM: Efficient Web Crawling for LLM Pretraining
Shi Yu12thanks: Work done while visiting Carnegie Mellon University.  Zhiyuan Liu1  Chenyan Xiong2 1Department of Computer Science and Technology, Tsinghua University 2Language Technologies Institute, Carnegie Mellon University yus21@mails.tsinghua.edu.cn; liuzy@tsinghua.edu.cn; cx@cs.cmu.edu
1 Introduction
Massive in size and diverse in topics, web data usually serve as the primary source of pretraining data for large language models (LLMs), providing an extensive and heterogeneous corpus that captures a wide spectrum of human knowledge and real-world information (Baack, 2024; Dubey et al., 2024; Penedo et al., 2024). Pretraining datasets are typically built from large-scale web crawls such as Common Crawl (CommonCrawl, 2007), which may contain TBs of data spanning billions of webpages (Penedo et al., 2024; Weber et al., 2024).
Figure 1: Graph traverse process of a traditional graph-connectivity-based crawler (green) and Craw4LLM (red) starting from a same seed URL (star).
Despite their vast scale, most of the data collected from web crawls are not used in the pretraining of LLMs. Existing work often discards over 90% of the raw data collected from the web (Li et al., 2024; Penedo et al., 2024; Tang et al., 2024), highlighting the inefficiency of current web crawlers in collecting LLM pretraining data. Common web crawlers like Common Crawl prioritize pages based on graph connectivity metrics like PageRank (Page et al., 1999; Cho et al., 1998) or harmonic centrality (Boldi and Vigna, 2014; Baack, 2024), which favor documents with a high number of inlinks (indegree) (Fortunato et al., 2008) rather than those most relevant for pretraining. This misalignment not only leads to waste in computational resources during excessive data processing for LLM developers, but also incentivizes over-crawling, which burdens website operators with redundant traffic and increases ethical and legal risks related to fair use of data and copyright (Longpre et al., 2024; New York Times, 2023).
To bridge this gap, we propose Web Crawling for LLM Pretraining (Craw4LLM). Instead of relying on traditional graph-connectivity-based signals, Craw4LLM improves crawling efficiency by prioritizing webpages based on their influence on LLM pretraining. Specifically, during each crawling iteration, all newly discovered documents are scored with a pretraining influence scorer derived from data-filtering pipelines for pretraining (Li et al., 2024; Penedo et al., 2024), and documents with the highest scores are used to discover new documents. By prioritizing webpages with high influence scores, as illustrated in Figure 1, Craw4LLM explores the web graph in a fundamentally different manner from traditional graph-connectivity-based crawlers, uncovering a distinct subset of the web more useful for pretraining.
We conduct large-scale crawling simulations on ClueWeb22-A (Overwijk et al., 2022), a snapshot of the web containing 900 million English webpages obtained from the central index of a commercial search engine. Results show that, by crawling only 1× of the pretraining dataset size, Craw4LLM can outperform traditional crawlers which collect 1×, 2×, and 4× data followed by data selection. Compared to the baseline crawler that achieves the same performance, Craw4LLM crawls only 21% of the webpages. Further analysis reveals that during crawling, Craw4LLM quickly discovers documents that align with the oracle selection, which selects from the full web graph. As a result, it achieves over 95% of the oracle performance while crawling only 2.2% of the data.
2 Methodology
(a) Pretraining (-0.11).
(b) PageRank (0.88).
Figure 2: Correlations between pretraining influence scores from DCLM fastText (Li et al., 2024) and PageRank to indegrees, on randomly sampled ClueWeb22-B documents (Overwijk et al., 2022). Spearman correlation coefficients are reported in parentheses.
In this section, we introduce Web Data Crawling for LLM Pretraining (Craw4LLM), an efficient crawling method that integrates LLM pretraining preference into the crawler. The algorithm of Craw4LLM is presented in Algorithm 1.
Algorithm 1 Craw4LLM Algorithm
0:  Seed URLs $\mathcal{U}_{\mathrm{seed}}$, number of pages to be crawled $N$, number of pages to be crawled in each iteration $n$, pretraining influence scorer $\mathcal{M}(\cdot)$
0:  Crawled page set $\mathcal{P}$
1:  Initialize URL and score priority queue $\mathcal{Q}\leftarrow\emptyset$
2:  Initialize crawled page set $\mathcal{P}\leftarrow\emptyset$
3:  Initialize visited URL set $\mathcal{V}\leftarrow\mathcal{U}_{\mathrm{seed}}$
4:  $\mathcal{U}_{c}\leftarrow\mathcal{U}_{\mathrm{seed}}$
5:  while $\lvert\mathcal{P}\rvert\leq N$ do
6:     $\mathcal{P}_{c}\leftarrow\textsc{FetchPages}(\mathcal{U}_{c})$
7:     Merge $\mathcal{P}_{c}$ into $\mathcal{P}$
8:     $\mathcal{U}_{\mathrm{out}}\leftarrow\textsc{ExtractURLs}(\mathcal{P}_{c})$
9:     for all $v\in\mathcal{U}_{\mathrm{out}}$ do
10:        if $v\notin\mathcal{V}$ then
11:           $\textsc{Enqueue}(\mathcal{Q},\;v,\;\textsc{Score\_URL}(v;\mathcal{M}))$
12:           $\textsc{Add}(\mathcal{V},\;v)$
13:        end if
14:     end for
15:     $\mathcal{U}_{c}\leftarrow\textsc{Dequeue}(\mathcal{Q},\;n)$
16:  end while
17:  return  $\mathcal{P}$
[…]
$s\leftarrow\textsc{Score\_URL}(u;\mathcal{M})=\mathcal{M}(\textsc{FetchPage}(u)),$ (1)
where $\textsc{FetchPage}(u)$ gets the page content of $u$ and $\mathcal{M}(\cdot)$ returns the score. Once all outlinks have been scored, following the standard procedures of existing crawlers, they are inserted into a priority queue, which automatically orders them based on their scores. The top $n$ highest-scoring URLs are then dequeued for pretraining and serve as the sources for the next round of crawling. This process repeats until $N$ documents have been collected, forming the final pretraining dataset $\mathcal{P}$.
Selection Commonsense Reasoning Language Understanding Reading Comprehension Symbolic Problem Solving World Knowledge Core % of
Method Pool Size (4 tasks) (6 tasks) (3 tasks) (5 tasks) (5 tasks) (23 tasks) Oracle
Using the DCLM fastText (Li et al., 2024) classifier as the pretraining influence scorer
Oracle 45× 0.2438 0.2209 0.1483 0.2039 0.2403 0.2239 100%
[…]
0.1896 0.1967 0.1260 0.2000 0.2024 0.1964 87.7%
Indegree 0.1730 0.1680 0.0326 0.1616 0.1668 0.1556 69.5%
0.1845 0.1856 0.0970 0.1958 0.1953 0.1865 83.3%
[…]
Using the FineWeb-Edu (Penedo et al., 2024) classifier as the pretraining influence scorer
Oracle 45× 0.1899 0.1973 0.1081 0.2117 0.2786 0.2133 100%
[…]
0.1797 0.1888 0.0931 0.1586 0.2100 0.1807 84.7%
Indegree 0.1730 0.1680 0.0326 0.1616 0.1668 0.1556 72.9%
0.1720 0.1709 0.0840 0.1783 0.1842 0.1724 80.8%
[…]
Table 1: Downstream LLM performance. Either the DCLM fastText classifier (Li et al., 2024) (top) or FineWeb-Edu classifier (Penedo et al., 2024) (down) is used as the pretraining influence scorer for Craw4LLM, and to select documents for oracle and crawl-then-select runs (Section 3). All models are pretrained on 1× data (20M documents, 32.9B tokens). The evaluation metric is centered accuracy (0 = random guess) (Li et al., 2024). Best/2nd best in the last two groups are bolded/underlined. See Appendix D for detailed results.
In contrast, traditional crawlers typically rely on graph connectivity metrics, such as PageRank (Cho et al., 1998) and harmonic centrality (Baack, 2024), which basically assign higher priority to pages with higher indegrees (Fortunato et al., 2008). As shown in Figure 2(a), the indegrees of webpages exhibit a poor correlation with the scores assigned by the DCLM fastText classifier, a pretraining influence scorer for identifying high-quality pretraining data (Li et al., 2024). This confirms that graph connectivity-based crawlers are inefficient in crawling pretraining data.
By incorporating a pretraining influence scorer, Craw4LLM traverses the web graph in a way that prioritizes high-quality pretraining documents. This makes the crawling more efficient and enables the discovery of documents dramatically different with connectivity-based crawlers.
3 Experimental Methodology
In this section, we introduce our experimental setup, with details on the crawler implementation and LLM training provided in Appendix A and B.
Craw4LLM.
To run experiments in our limited computational budget, we run a simulation of Craw4LLM on the ClueWeb22 dataset (Overwijk et al., 2022), a snapshot of the web with graph information from a commercial crawler. We use the English subset of ClueWeb22-A, which is a web graph containing 900M webpages with links. We randomly sampled 10K URLs as our seed URLs. We set the number of total crawled documents $N$ to 20M and crawled documents each iteration $n$ to 10K. In our experiments, we consider using the DCLM fastText classifier (Li et al., 2024) or the FineWeb-Edu classifier (Penedo et al., 2024) as the pretraining influence scorer $\mathcal{M}(\cdot)$, whose details can be found in Appendix C.
Baselines.
We emulate traditional graph-connectivity-based crawlers by replacing the LLM-oriented URL scoring function (Eq. 1) with a function that returns the indegree for a given URL, since a node’s indegree closely correlates with PageRank, a common graph connectivity metric, as shown in Figure 2(b) and previous findings (Fortunato et al., 2008). We also introduce a random crawling baseline, where the scorer assigns random scores. We run both of them in a crawl-then-select setting, first crawling 1× or 2× documents and then selecting the top 1× (20M) documents based on scores assigned by the DCLM fastText classifier or the FineWeb-Edu classifier. This process mimics existing data-filtering pipelines, which begin with crawled documents and then apply filtering (Li et al., 2024; Penedo et al., 2024).
Oracle.
We introduce an oracle selection run in which we directly apply the pretraining influence scorer to the ClueWeb22-A dataset and randomly sample 20M documents which are scored as the 10% for pretraining, serving as the upper bound.
LLM Training and Evaluation.
For all runs, we use the final set of 20M crawled or selected documents to pretrain a 411M Transformer on 4× Chinchilla-optimal tokens (Hoffmann et al., 2022), totaling 32.9B tokens. The pretraining is conducted using the DCLM codebase (Li et al., 2024). To evaluate the pretrained LLMs, we follow the DCLM evaluation recipe, assessing performance on 23 (22 unique) core tasks.
4 Evaluation Results
In this section, we first present the overall performance of Craw4LLM (Section 4.1), followed by further analysis (Section 4.2).
4.1 Overall Performance
In this experiment, we compare the performance of Craw4LLM with baseline crawlers by evaluating LLMs trained on their respective crawled data. As shown in Table 1, when all methods crawl the same amount of training data (1×), Craw4LLM significantly outperforms random crawling and indegree crawling using either DCLM fastText or FineWeb-Edu classifier as the pretraining influence scorer. In the crawl-then-select setting, where traditional crawlers are allowed to collect twice as much data (2×) for later selection, they still underperform compared to Craw4LLM. This suggests that incorporating pretraining-oriented signals early in the crawling process is more beneficial than relying on post-selection. With only 1× of the data, Craw4LLM retains 95.3% and 95.6% of the performance achieved by the DCLM fastText and FineWeb-Edu oracle run, respectively, which directly selects from the entire ClueWeb22-A dataset, a substantially larger 45× data pool.
4.2 Analysis
In this subsection, we further analyze the efficiency of Craw4LLM compared to traditional crawlers and explore the reasons behind it. We utilize the DCLM fastText classifier in the experiments presented in this subsection.
(a) Extended crawling.
(b) Visited documents.
Figure 3: Efficiency of crawlers. (a) shows the performance of LLMs trained on selected data crawled by Craw4LLM and extended baseline crawlers. (b) presents the number of crawled ($\mathcal{P}$) and visited ($\mathcal{V}$) documents for Craw4LLM, along with the estimated number of crawled documents required for indegree-based crawler to match Craw4LLM’s performance.
Crawling Efficiency.
We evaluate the efficiency of Craw4LLM by comparing the number of documents it crawls or visits against baseline crawlers. As shown in Figure 3(a), even when the baselines crawl 4× the required pretraining data for selection, they still underperform compared to Craw4LLM. Extrapolation suggests that the indegree-based crawler would need to crawl 4.8× documents (96M) to match Craw4LLM’s performance. Figure 3(b) further illustrates that Craw4LLM achieves the same performance while crawling only 21% of the documents required by the indegree-based crawler, or 48% when considering all visited documents. These results highlight the efficiency of Craw4LLM, demonstrating its potential to reduce website burdens and mitigate over-crawling.
Document Coverage.
Figure 4: Precision (left) and recall (right) of the oracle documents among the documents crawled by Craw4LLM, indegree, and random crawler. The upper bound represents always crawling the oracle documents.
(a) 1-hop (0.61).
(b) 2-hop (0.60).
Figure 5: Correlations between the pretraining influence scores of the documents themselves and the average scores of their 1- and 2-hop outlink documents. Spearman correlation coefficients are reported in parentheses.
In this experiment, we plot the precision and recall of the oracle-selected documents among those crawled by Craw4LLM and baseline crawlers throughout the crawling process. As shown in Figure 4, the precision quickly reaches 1.0, while the recall increases linearly, aligning with the theoretical upper bound. The saturated performance remains until 13 million documents have been crawled, after which the performance starts to decline, likely due to the lack of connectivity of the ClueWeb22 subgraph. In contrast, baseline crawlers exhibit minimal overlap with oracle-selected data, verifying that most of their crawled content is misaligned with pretraining needs and should be filtered (Li et al., 2024; Penedo et al., 2024). These results emphasize the importance of targeted crawling strategies for pretraining.
Score Correlations Across Links.
Craw4LLM tracks the outlinks of the highest-scored documents in the current iteration to enrich the queue for future crawls. As shown in Figure 5, we plot the correlations between the pretraining influence scores of current documents and their 1- and 2-hop outlinks. The results indicate a correlation in influence scores across link hops, suggesting that highly-rated documents are interconnected and can be discovered through previously crawled documents.
5 Conclusion
This paper presents Craw4LLM, a step toward more efficient and responsible web crawling for LLM pretraining. By prioritizing documents based on the pretraining needs, our method improves crawling efficiency and reduces unnecessary crawling, easing the burden on web hosts. While fair use of web data remains a critical challenge, we hope that Craw4LLM can help mitigate these concerns and promote more compliant and sustainable practices in obtaining pretraining data for LLMs.
Limitations
Web crawling raises important concerns regarding copyright and the fair use of web data (Longpre et al., 2024), necessitating a better solution from the entire LLM community, such as sharing benefits with website owners. In this paper, we propose a more efficient crawling method that mitigates these challenges by reducing crawling, though it does not fully resolve them. Our experiments are conducted on a web graph dataset ClueWeb22 (Overwijk et al., 2022), thereby avoiding issues associated with actual web crawling. We hope that future advancements in web crawling will better align with ethical and legal standards.
While our crawling simulation is a sufficient research setup, further validation is required to assess the effectiveness of Craw4LLM in real-world crawling scenarios. Our Craw4LLM and baseline crawlers implement only the selection policy (Cho et al., 1998) of a crawler, which determines which pages to crawl. Although we try to mimic real-world crawling procedures used in systems like Apache Nutch [^1] https://nutch.apache.org/, we do not implement other web crawling policies in industrial-level crawlers, such as the re-visit policy (Cho and Garcia-Molina, 2003a), politeness policy (Cho and Garcia-Molina, 2003b), and parallelization policy (Cho and Garcia-Molina, 2002). We leave the integration of Craw4LLM into real-world crawling engines like Nutch and a comprehensive comparison between Craw4LLM and traditional crawling methods in real-world crawling scenarios for future work.
Appendix A Details on Crawling
Our implementation of the indegree-based crawler employs a static URL scoring function, which directly returns the indegree of a given URL based on the full ClueWeb22 graph (Sec. 3). For real-world crawlers, as the true indegree value of a URL cannot be known in advance, a local graph must be maintained to track the known inlinks of discovered URLs. This local graph is updated iteratively as the discovered portion of the web expands during the crawling process (Cho et al., 1998).
Maintaining such a local graph during crawling introduces significant computational overhead. For simplicity, we instead implement the static simulation, where we directly return the global indegree of each URL. We believe that this simplified implementation does not underperform compared to real-world implementations, as our approach leverages global information from the entire graph, which should be better than the partial information from the local graph.
We run our simulated crawlers on a Linux server equipped with two Intel(R) Xeon(R) E5-2630 v3 CPUs (8 cores per socket, 16 cores in total, 1 thread per core), 125GiB of memory, and an SSD. The crawling process of Craw4LLM takes about 21 hours to finish. In comparison, the random and indegree-based crawlers take around 10.5 and 12.5 hours, respectively. Note that since these are simulated crawls on the snapshot, the reported times do not reflect real-world crawling performance.
Appendix B Details on LLM Training and Evaluation
Hyper-parameter Value
$n_{\text{layers}}$ 24
$n_{\text{heads}}$ 8
$d_{\text{model}}$ 1,024
$d_{\text{head}}$ 128
Warmup 2,000
Learning Rate 3e-3
Weight Decay 0.033
z-loss 1e-4
Global Batch Size 512
Sequence Length 2048
Table 2: Model and training hyper-parameters. $n_{\text{layers}}$, $n_{\text{layers}}$, $d_{\text{model}}$, and $d_{\text{head}}$ denote the number of layers, attention heads, width, and width per attention head, respectively.
We pretrain a 411M-parameter [^2] Sometimes referred to as 400M in the DCLM paper (Li et al., 2024). decoder-only Transformer model using the DCLM training recipe (Li et al., 2024) [^3] https://github.com/mlfoundations/dclm. The hyper-parameters are presented in Tabel 2. To enhance training stability, we extend the original 411M-1x setting to 411M-4x, meaning the model is trained on 4 times the Chinchilla-optimal number of tokens (Hoffmann et al., 2022), which amounts to 32.9B tokens. The training process takes 1 day and 12 hours on 8 NVIDIA L40S GPUs. For further details, please refer to the DCLM paper (Li et al., 2024). Due to computational constraints, each pretraining experiment is conducted only once.
We use the DCLM evaluation recipe (Li et al., 2024) to evaluate model performance on 23 (22 unique) core tasks.
Appendix C Details on the Pretraining Influence Scorers
DCLM fastText Classifier
(Li et al., 2024) is a quality filter based on fastText (Joulin et al., 2017), trained to distinguish between high- and low-quality pretraining data. A dataset size of 400K examples (200K positive, 200K negative) is used for training. Negative samples are random documents from an earlier version of the RefinedWeb reproduction from the DCLM team. Positive samples come from OpenHermes 2.5 (Teknium, 2023) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. The final DCLM-baseline dataset is created by selecting only the top 10% of documents ranked by the classifier.
FineWeb-Edu Classifier
(Penedo et al., 2024) is a model designed to identify and filter educational content within FineWeb. Based on the Snowflake-arctic-embed-m [^4] https://huggingface.co/Snowflake/snowflake-arctic-embed-m text embedding model, it is trained on 450,000 annotations generated by Llama-3-70B-Instruct, which assigned an educational quality score from 0 to 5 to every document. FineWeb-Edu is built by filtering out samples from FineWeb with scores lower than 3 with the trained classifier.
Appendix D Detailed Results
Pretraining Selection Commonsense Reasoning
Method Influence Scorer Pool Size CommonsenseQA COPA OpenBookQA PIQA
Oracle DCLM fastText 45× 0.2850 0.7000 0.3300 0.6812
Oracle FineWeb-Edu 45× 0.2342 0.6200 0.3620 0.6638
Random 0.2072 0.6700 0.2980 0.6746
Random DCLM fastText 0.2588 0.6200 0.3160 0.6785
Random FineWeb-Edu 0.2301 0.6200 0.3140 0.6779
Random DCLM fastText 0.2326 0.6400 0.3380 0.6757
Indegree 0.3219 0.6000 0.2780 0.6513
Indegree DCLM fastText 0.1966 0.6600 0.3040 0.6752
Indegree FineWeb-Edu 0.1974 0.6400 0.3160 0.6616
Indegree DCLM fastText 0.2088 0.6400 0.3400 0.6817
Craw4LLM DCLM fastText 0.2277 0.6600 0.3300 0.6926
Craw4LLM FineWeb-Edu 0.3219 0.6200 0.3500 0.6616
Table 3: Results for commonsense reasoning tasks.
Pretraining Selection Language Understanding
Method Influence Scorer Pool Size BIG-Bench Lang. Id. HellaSwag (zero-shot) HellaSwag LAMBADA Winograd Winogrande
Oracle DCLM fastText 45× 0.2515 0.3856 0.3905 0.4432 0.6557 0.5130
Oracle FineWeb-Edu 45× 0.2580 0.3831 0.3818 0.3643 0.6227 0.5185
Random 0.2490 0.3709 0.3716 0.3990 0.6044 0.5146
Random DCLM fastText 0.2468 0.3882 0.3925 0.4073 0.6007 0.5130
Random FineWeb-Edu 0.2485 0.3815 0.3800 0.3804 0.6007 0.5146
Random DCLM fastText 0.2521 0.4011 0.4019 0.4390 0.6154 0.5130
Indegree 0.2566 0.3515 0.3519 0.3596 0.5971 0.5004
Indegree DCLM fastText 0.2547 0.3749 0.3771 0.3773 0.5861 0.5241
Indegree FineWeb-Edu 0.2528 0.3636 0.3658 0.3672 0.5678 0.5193
Indegree DCLM fastText 0.2562 0.3994 0.4008 0.4159 0.6190 0.5178
Craw4LLM DCLM fastText 0.2544 0.4035 0.4048 0.4196 0.6593 0.5288
Craw4LLM FineWeb-Edu 0.2521 0.3726 0.3717 0.3478 0.6264 0.5083
Table 4: Results for language understanding tasks.
Pretraining Selection Reading Comprehension
Method Influence Scorer Pool Size BoolQ CoQA SQuAD
Oracle DCLM fastText 45× 0.5755 0.2479 0.3139
Oracle FineWeb-Edu 45× 0.5367 0.2305 0.3130
Random 0.5080 0.1799 0.1882
Random DCLM fastText 0.5807 0.2053 0.2759
Random FineWeb-Edu 0.5627 0.1997 0.2304
Random DCLM fastText 0.5911 0.2361 0.2951
Indegree 0.5324 0.1666 0.1616
Indegree DCLM fastText 0.5697 0.1843 0.2390
Indegree FineWeb-Edu 0.5670 0.1904 0.2011
Indegree DCLM fastText 0.5765 0.2147 0.2736
Craw4LLM DCLM fastText 0.5440 0.2264 0.2215
Craw4LLM FineWeb-Edu 0.4654 0.2146 0.3026
Table 5: Results for reading comprehension tasks.
Pretraining Selection Symbolic Problem Solving
Method Influence Scorer Pool Size AGI Eval LSAT-AR BIG-Bench CS Algorithms BIG-Bench Dyck Lang. BIG-Bench Operators BIG-Bench Repeat Copy Logic
Oracle DCLM fastText 45× 0.2739 0.4341 0.2160 0.2143 0.0625
Oracle FineWeb-Edu 45× 0.2826 0.4606 0.2870 0.1762 0.0313
Random 0.2391 0.4568 0.1970 0.2143 0.0000
Random DCLM fastText 0.2696 0.4538 0.2520 0.1762 0.0313
Random FineWeb-Edu 0.2000 0.4091 0.2050 0.1476 0.0313
Random DCLM fastText 0.1957 0.4568 0.2600 0.1857 0.0625
Indegree 0.2304 0.4371 0.1900 0.1429 0.0000
Indegree DCLM fastText 0.2609 0.4235 0.2340 0.2143 0.0313
Indegree FineWeb-Edu 0.2304 0.4545 0.2060 0.1619 0.0313
Indegree DCLM fastText 0.2174 0.4538 0.2530 0.1667 0.0938
Craw4LLM DCLM fastText 0.2696 0.4371 0.1620 0.2095 0.0938
Craw4LLM FineWeb-Edu 0.2609 0.4750 0.1780 0.2048 0.0938
Table 6: Results for symbolic problem solving tasks.
Pretraining Selection World Knowledge
Method Influence Scorer Pool Size ARC Easy ARC Challenge BIG-Bench QA Wikidata Jeopardy MMLU
Oracle DCLM fastText 45× 0.5951 0.3166 0.4945 0.1176 0.2805
Oracle FineWeb-Edu 45× 0.6343 0.3549 0.5407 0.1726 0.2706
Random 0.5152 0.2799 0.5186 0.0461 0.2552
Random DCLM fastText 0.5425 0.2807 0.5081 0.0648 0.2561
Random FineWeb-Edu 0.5535 0.2816 0.5194 0.0863 0.2483
Random DCLM fastText 0.5577 0.2867 0.5126 0.0970 0.2543
Indegree 0.4857 0.2509 0.4888 0.0138 0.2618
Indegree DCLM fastText 0.5248 0.2790 0.5205 0.0555 0.2464
Indegree FineWeb-Edu 0.5143 0.2679 0.5044 0.0553 0.2389
Indegree DCLM fastText 0.5749 0.2935 0.5084 0.0959 0.2430
Craw4LLM DCLM fastText 0.6103 0.3208 0.5143 0.1323 0.2661
Craw4LLM FineWeb-Edu 0.6427 0.3592 0.5319 0.1859 0.2736
Table 7: Results for world knowledge tasks.
The raw (uncentered) accuracy of all evaluation tasks is presented in Table 3, 4, 5, 6, and 7. Please refer to Li et al. (2024) for more details on the evaluation tasks.
Appendix E The ClueWeb22 Dataset
ClueWeb22 (Overwijk et al., 2022) is distributed under a “TREC-style” license for research purpose. The dataset can be obtained by signing a data license agreement with Carnegie Mellon University [^5] https://lemurproject.org/clueweb22/obtain.php. We use ClueWeb22 only for research purpose.
Appendix F Use of AI Assistants
We use GitHub Copilot [^6] https://github.com/features/copilot to assist with coding and ChatGPT [^7] https://chatgpt.com/ (powered by GPT-4o) to enhance the writing of this paper.